# Improved Bounds for the Freiman-Ruzsa Theorem

### v1

@@ 1 Introduction
It turns out that any set of small doubling can be “built” out of these two examples. In the case $G=\mathbb{Z}^{d}$, this was first proved by Freiman ([freiman1973foundations], [freiman1966nachala]), in the case of $G$ with bounded torsion by Ruzsa ([RuzsaFiniteTorsion]), and for a general abelian group $G$, by Green and Ruzsa [GRFreiman]:
@@
The three appearances of $f(K)$ written above can be considered three different functions, and in some cases it is possible to decrease one of the bounds at the cost of increasing another. We will not need to do this, but a description of these maneuvers (and their tradeoffs) can be found in [SRevisited, Section 12].
The argument of Green and Ruzsa in [GRFreiman] shows that Question 1.2 is true for $f(K)=O(K^{4}\log^{O(1)}(2K))$. There have since been many improvements to these bounds. In [sanders2012bogolyubov], Sanders showed that one can take $f(K)=\log^{O(1)}(2K)$, and combining this with a bootstrapping argument by Konyagin, he established the best-known bound:
@@
Question 1.2 is true for a function $f(K)$ with [^2] The constant 16 is present to ensure that $f(K)$ is positive even when $K$ is very close to 1, and can otherwise be disregarded.
$f(K)=O(\log(16K)\log\log(16K)\log\log\log(16K)).$
@@
Conjecture 1.5 has a number of equivalent formulations; for example, it is equivalent to the corresponding statement where the set of all convex bodies is replaced with the set of ellipsoids [Mannersformulations]. Another possible variation, where the set of all convex bodies is replaced with the set of all axis-aligned boxes, is false and was disproved by Lovett and Regev [Lovett2017counterexample].
In the case where $G$ has bounded torsion, Conjecture 1.5 was originally formulated by Marton, and proved by Gowers, Green, Manners, and Tao in [GGMTBoundTorsion] and [GGMTTwoTorsion]. The main result from [GGMTBoundTorsion] is:
@@
Comparing this with Theorem 1.4, we see that our bound improves on theirs in the regime where $m^{3}$ is much larger than $\log\log(K)\log\log\log(K)$.
One may also consider the variant of Question 1.2 where the restriction on the size of the convex coset progression is lifted, so that one only cares about covering $A$ efficiently by translates of a low-dimensional set. Showing that $A$ can be covered by at most $\exp(O(\log(2K))$ translates of a convex coset progression with dimension at most $O(\log(2K))$ is sometimes called the weak Polynomial Freiman-Ruzsa conjecture. In [GMTRevisited], it was shown that Marton’s conjecture in $\mathbb{F}_{2}^{n}$ implies the weak Polynomial Freiman-Ruzsa conjecture, so this problem was solved in [GGMTTwoTorsion].
@@ 1.1 Related Problems and Applications
A very useful alternative structural result for sets of small doubling is one that provides a large structured set, such as a convex coset progression, contained in A+AAAA+A-A-A. A result of this kind, for dense subsets of a finite group, was first proved by Bogolyubov [bogoliouboff1939quelques] and was extended to arbitrary sets with small doubling by Ruzsa [ruzsa1994generalized] en route to a proof of Freiman’s theorem. The best known result in this direction is due to Sanders [sanders2012bogolyubov]:
@@
In most previous proofs of Freiman-Ruzsa type theorems, a Bogolyubov-Ruzsa type result is proved as an intermediate step. This is not the case for the arguments in [GGMTBoundTorsion] and [GGMTTwoTorsion], whose arguments only find a structured set contained in a sum of $\log^{O_{m}(1)}(2K)$ copies of $A$. Our methods perform even more poorly at this task, and do not directly prove that there is a large convex coset progression inside $\underbrace{A+A+\dots+A}_{n\text{ times}}$ for any $n$.
We now turn our attention towards applications of Theorem 1.4. One important application of structural results for sets with small doubling is an inverse theorem for the Gowers $U^{3}(G)$ norm, as proved by Green and Tao in [GTinverse]. There, Green and Tao apply the Bogolyubov-Ruzsa lemma to prove the $U^{3}(G)$ inverse theorem. However, Samorodnitsky [Samorodnitsky] showed that for $G=\mathbb{F}_{2}^{n}$, Marton’s conjecture is sufficient to obtain polynomial bounds for the $U^{3}(G)$ inverse theorem. A similar implication when $G=\mathbb{F}_{p}^{n}$ for an odd prime $p$ was proved by Gowers, Green, Manners, and Tao in [GGMTBoundTorsion, Appendix C].
One may also modify the arguments in [GTinverse] to obtain improved bounds on the dimension in the $U^{3}$ inverse theorem.
@@
We sketch a proof of Theorem 1.8 in Appendix C. We note that the $\log(\eta^{-1})^{1+o(1)}$ terms can be replaced by $\log(\eta^{-1})\log\log^{O(1)}(\eta^{-1})$ if desired. However, we do not obtain a bound similar in shape to that in Theorem 1.4, i.e., $\log(\eta^{-1})\log\log(\eta^{-1})\log\log\log(\eta^{-1})$.
There are a number of other applications of Freiman-Ruzsa-type bounds, many mentioned in [SRevisited], such as $\Lambda(4)$ estimates for the squares and lower bounds for cardinalities $|A+\alpha A|$, where $A\subseteq\mathbb{R}$ is finite and $\alpha$ is transcendental. Our results provide no new information other than an improved input for these existing arguments, so we invite the reader to explore [SRevisited, Section 13] and the references therein for details.
Finally, as an application of Marton’s conjecture, Prendiville [Prendiville] proved an efficient inverse theorem for the Gowers $U^{3}(\mathbb{F}_{p}^{n})$ norm relative to quadratic level sets, and then gave an application to estimating the Ramsey number of Brauer quadruples. The author would be interested to see if a similar result is possible for general finite abelian groups $G$, and whether bounds of the quality in Theorem 1.4 can lead to good bounds for combinatorial configuration problems like those considered in [Prendiville].
@@ Acknowledgements
The author would like to thank Terence Tao for a great deal of guidance and support. The author would also like to thank James Leng for helpful conversations and encouragement, especially regarding the $U^{3}$ inverse theorem. The author is partially supported by NSF Grant DMS-2347850.
@@ 2.1 Definitions and Notation
To simplify notation in this paper, we introduce the following conventions regarding iterated logarithms.
@@
For a real number $x\geq 16$, we define $\log_{2}(x)=\log\log(x)$ and $\log_{3}(x)=\log\log\log(x)$. In particular, we never use $\log_{b}$ to refer to the base $b$-logarithm.
@@
Entropy analogues of standard inequalities in additive combinatorics were developed by Ruzsa [Ruzsaentropy] and Tao [TSumsetEntropy] in 2009. Several authors have contributed to this program since then, developing an “entropic Ruzsa calculus” which allows us to perform many of the standard operations from additive combinatorics in the entropic setting. We record the estimates from this theory which we will use in Appendix A.
@@
Although the doubling constant and Ruzsa distance provide important information about the growth and dimension of a set $A\subseteq G$, they do not paint a complete picture of the approximate structure of $A$. We consider the following pair of standard examples, also discussed in [SRevisited] and [GGMTTwoTorsion].
@@
To distinguish between these instances of “covering” and “containment”, we define the following, which is analogous to a combinatorial polynomial growth condition studied by Sanders in [sanders2012bogolyubov], [SRevisited].
@@ 2.2 Strategy of Proof
Part 1: The Entropy Side. To motivate the first half of our argument, we will give a brief overview of the proof of the Polynomial Freiman-Ruzsa conjecture over $\mathbb{F}_{2}$ given by Gowers, Green, Manners, and Tao [GGMTTwoTorsion]. As in [GGMTTwoTorsion], we will obtain a structural characterization of the minimizers of a functional involving some entropic Ruzsa distances. More precisely, they prove the following.
Theorem 2.9 ([GGMTTwoTorsion], Propositions 2.1 and 2.2).
@@
Let $\eta=1/1009$. For random variables $X,Y$ supported on $A$, we define
$\tau[X;Y]=\text{d}[X;Y]+\frac{\eta}{\log_{2}(16K)\log_{3}(16K)}(\text{d}[X;U_{A}]+\text{d}[Y;U_{A}]).$
@@
Let $C$ be a constant. Suppose $(X,Y)$ are minimizers of $\tau$. Then either $K=O_{C}(1)$, or $X$ has polynomial growth of order $d=O(\log(16K)\log_{3}(16K))$ at scale $Cd^{5}$. Moreover, $\text{d}[X;U_{A}]\leq 30d\log(d)$.
The constants 1/1009 and 30 above can certainly be improved, but we make no effort to do so.
Part 2: The Fourier Side. The arguments in this part of the proof largely mirror those by Sanders in [SRevisited, Sections 9-11]. Sanders considered a combinatorial polynomial growth condition instead of an entropic one, and showed that a set with polynomial growth of a given order could be related to a convex coset progression with controlled size and dimension. More specifically, he showed the following.
Theorem 2.12 ([SRevisited], Theorem 2.7).
@@
The condition (2) is stronger than Definition 2.8 in the sense that if a set $X$ satisfies condition (2) with an exponent $d$, then $U_{X}$ has polynomial growth of order $d$ as well. The converse implication is not true; see [GMTRevisited, Appendix B] for an example of large gaps between combinatorial and entropic notions of doubling. However, we are able to show using modifications of Sanders’ arguments that under our weaker notion of polynomial growth, we may obtain a comparable amount of structure.
@@
In view of the quantitative similarities between Theorem 2.12 and Proposition 2.13, it is natural to wonder how our argument provides any quantitative improvement over the results in [SRevisited]. There, Sanders uses the Croot-Sisask lemma for almost-periods of convolutions [CrootSisaskAP] to produce a somewhat sparse set $T$ ($\text{d}[U_{T};U_{A}]\approx\log(K)^{3}$) with a relatively high order of polynomial growth ($d\approx\log(K)^{3}$). The information-theoretic part of our argument produces a much “denser” random variable with a much lower order of polynomial growth, which eventually leads to savings in the final bounds. The author would be very interested to see if the methods in this paper can be used to provide quantitative improvements to any of the many other applications of the Croot-Sisask lemma.
@@
Let $C$ be the constant from Proposition 2.13. We now select a minimizer $(X,Y)$ of $\tau$ and apply Proposition 2.11. If $K=O(1)$, we are done, since we can apply Theorem 1.3 to cover $A$ by $O(1)$ translates of a convex coset progression of size at most $O(|A|)$ and dimension at most $O(1)$. Otherwise, $X$ and $A$ satisfy the hypotheses of Proposition 2.13 with $d=O(\log(16K)\log_{3}(16K))$. There is thus a convex coset progression $M$ with dimension $O(d\log(d))$ and size at most $\exp(O(d\log(d))|A|$ such that $A$ can be covered by at most $K\exp(O(d\log(d)))$ translates of $M$. Since $d\log(d)=O(\log(16K)\log_{2}(16K)\log_{3}(16K))$, the claim follows. ∎
@@ 3 Obtaining a Random Variable with Polynomial Growth
$\text{d}[X|nX;Y|nY]\leq\log(k)-\log(k)(\log_{2}(16n)-\log_{2}(16(n-1))\leq\log(k)-\frac{\log(k)}{n\log(16n)}.$
Provided $K\geq 16$, $7\log_{2}(16K)\geq\log(16\log(K)^{6})\geq\log(16n)$, so the above quantity is at most
$\log(k)-\frac{\log(k)}{7n\log_{2}(16K)}.$
@@
$\frac{72}{n}\log(k)\log_{3}(16K)$
since $m\geq n/2$ and $m\leq\log(K)^{6}$.
By a similar argument, $\text{d}[Y|nY;U_{A}]-\text{d}[Y;U_{A}]\leq\frac{72}{n}\log(k)\log_{3}(16K)$.
@@
$\text{d}[X|nX;Y|nY]-\text{d}[X;Y]\leq-\frac{\log(k)}{7n\log_{2}(16K)},$
$\frac{\eta}{\log_{2}(16K)\log_{3}(16K)}(\text{d}[X|nX;U_{A}]-\text{d}[X;U_{A}])\leq\frac{72\eta\log(k)}{n\log_{2}(16K)},\text{ and}$
$\frac{\eta}{\log_{2}(16K)\log_{3}(16K)}(\text{d}[Y|nY;U_{A}]-\text{d}[Y;U_{A}])\leq\frac{72\eta\log(k)}{n\log_{2}(16K)}.$
@@
$\sum_{s,t\in G}\mathbb{P}(nX=t,nY=s)\tau[X|nX=t;Y|nY=s]-\tau[X;Y]\leq-\frac{\log(k)}{7n\log_{2}(16K)}+\frac{144\eta\log(k)}{n\log_{2}(16K)},$
and the right-hand side is strictly negative since $\eta<\frac{1}{7\cdot 144}$. There thus exist $s,t\in G$ such that $\tau[X|nX=t;Y|nY=s]<\tau[X;Y]$, so $(X,Y)$ could not have been a minimizer of $\tau$. ∎
@@
By Proposition 3.1, either $K=O(1)$, or this quantity is bounded above by
$6\sum_{j=1}^{\ell}\log(k)+\log(k)\log_{2}(2^{j})\leq 6\ell\log(k)+6\log(6)\ell\log(k)+6\ell\log_{3}(K)\leq 24\ell\log_{3}(16K).$
@@
$\mathbb{H}(nX)-\mathbb{H}(X)\leq\mathbb{H}(2^{\ell}X)-\mathbb{H}(X)\leq 24\ell\log(k)\log_{3}(16K)\leq 48\log(n)\log(k)\log_{3}(16K).$
@@
$\log(k)=\text{d}[X;Y]\leq\tau[X;Y]\leq\tau[U_{A};U_{A}]\leq 3\log(K)\leq 3\log(16K).$
Thus for any $n\leq\frac{1}{4}\log(K)^{6}$,
$\mathbb{H}(nX)-\mathbb{H}(X)\leq 144\log(n)\log(16K)\log_{3}(16K).$
Set $d=144\log(16K)\log_{3}(16K)$. Since $d^{5}=o(\log(K)^{6})$, either $K=O_{C}(1)$, or
@@
It remains to show that $\text{d}[X;U_{A}]\leq Cd\log(d)$. Since $(X,Y)$ minimizes $\tau$, we have
$\frac{\eta}{\log_{2}(16K)\log_{3}(16K)}\text{d}[X;U_{A}]\leq\tau[X;Y]\leq\tau[U_{A};U_{A}]\leq 3\text{d}[U_{A};U_{A}]\leq 3\log(K),$
@@
$\text{d}[X;U_{A}]\leq 3027\log(K)\log_{2}(16K)\log_{3}(16K).$
Either $K=O(1)$, or this is at most $30d\log(d)$. ∎

### v2

@@ 1 Introduction
It turns out that any set of small doubling can be “built” out of these two examples. In the case $G=\mathbb{Z}^{d}$, this was first proved by Freiman ([5], [6]), in the case of $G$ with bounded torsion by Ruzsa ([24]), and for a general abelian group $G$, by Green and Ruzsa [12]:
@@
The three appearances of $f(K)$ written above can be considered three different functions, and in some cases it is possible to decrease one of the bounds at the cost of increasing another. We will not need to do this, but a description of these maneuvers (and their tradeoffs) can be found in [27, Section 12].
The argument of Green and Ruzsa in [12] shows that Question 1.2 is true for $f(K)=O(K^{4}\log^{O(1)}(2K))$. There have since been many improvements to these bounds. In [26], Sanders showed that one can take $f(K)=\log^{O(1)}(2K)$, and combining this with a bootstrapping argument by Konyagin, he established the best-known bound:
@@
Question 1.2 is true for a function $f(K)$ with [^2] The constant 3 is present to ensure that $f(K)$ is positive even when $K$ is very close to 1, and can otherwise be disregarded.
$f(K)=O(\log(3K)\log\log(3K)).$
@@
Conjecture 1.5 has a number of equivalent formulations; for example, it is equivalent to the corresponding statement where the set of all convex bodies is replaced with the set of ellipsoids [19]. Another possible variation, where the set of all convex bodies is replaced with the set of all axis-aligned boxes, is false and was disproved by Lovett and Regev [14].
In the case where $G$ has bounded torsion, Conjecture 1.5 was originally formulated by Marton, and proved by Gowers, Green, Manners, and Tao in [9] and [10]. The main result from [9] is:
@@
Comparing this with Theorem 1.4, we see that our bound improves on theirs in the regime where $m^{3}$ is much larger than $\log\log(K)$.
One may also consider the variant of Question 1.2 where the restriction on the size of the convex coset progression is lifted, so that one only cares about covering $A$ efficiently by translates of a low-dimensional set. Showing that $A$ can be covered by at most $\exp(O(\log(2K))$ translates of a convex coset progression with dimension at most $O(\log(2K))$ is sometimes called the weak Polynomial Freiman-Ruzsa conjecture. In [11], it was shown that Marton’s conjecture in $\mathbb{F}_{2}^{n}$ implies the weak Polynomial Freiman-Ruzsa conjecture, so this problem was solved in [10].
@@ 1.1 Related Problems and Applications
A very useful alternative structural result for sets of small doubling is one that provides a large structured set, such as a convex coset progression, contained in A+AAAA+A-A-A. A result of this kind, for dense subsets of a finite group, was first proved by Bogolyubov [1] and was extended to arbitrary sets with small doubling by Ruzsa [23] en route to a proof of Freiman’s theorem. The best known result in this direction is due to Sanders [26]:
@@
In most previous proofs of Freiman-Ruzsa type theorems, a Bogolyubov-Ruzsa type result is proved as an intermediate step. This is not the case for the arguments in [9] and [10], whose arguments only find a structured set contained in a sum of $\log^{O_{m}(1)}(2K)$ copies of $A$. Our methods perform even more poorly at this task, and do not directly prove that there is a large convex coset progression inside $\underbrace{A+A+\dots+A}_{n\text{ times}}$ for any $n$.
We now turn our attention towards applications of Theorem 1.4. One important application of structural results for sets with small doubling is an inverse theorem for the Gowers $U^{3}(G)$ norm, as proved by Green and Tao in [13]. There, Green and Tao apply the Bogolyubov-Ruzsa lemma to prove the $U^{3}(G)$ inverse theorem. However, Samorodnitsky [25] showed that for $G=\mathbb{F}_{2}^{n}$, Marton’s conjecture is sufficient to obtain polynomial bounds for the $U^{3}(G)$ inverse theorem. A similar implication when $G=\mathbb{F}_{p}^{n}$ for an odd prime $p$ was proved by Gowers, Green, Manners, and Tao in [9, Appendix C].
One may also modify the arguments in [13] to obtain improved bounds on the dimension in the $U^{3}$ inverse theorem.
@@
We sketch a proof of Theorem 1.8 in Appendix C. We note that the $\log(\eta^{-1})^{1+o(1)}$ terms can be replaced by $\log(\eta^{-1})\log\log^{O(1)}(\eta^{-1})$ if desired. However, we do not obtain a bound of exactly the same shape as that in Theorem 1.4, i.e., $\log(\eta^{-1})\log\log(\eta^{-1})$.
There are a number of other applications of Freiman-Ruzsa-type bounds, many mentioned in [27], such as $\Lambda(4)$ estimates for the squares and lower bounds for cardinalities $|A+\alpha A|$, where $A\subseteq\mathbb{R}$ is finite and $\alpha$ is transcendental. Our results provide no new information other than an improved input for these existing arguments, so we invite the reader to explore [27, Section 13] and the references therein for details.
Finally, as an application of Marton’s conjecture, Prendiville [21] proved an efficient inverse theorem for the Gowers $U^{3}(\mathbb{F}_{p}^{n})$ norm relative to quadratic level sets, and then gave an application to estimating the Ramsey number of Brauer quadruples. The author would be interested to see if a similar result is possible for general finite abelian groups $G$, and whether bounds of the quality in Theorem 1.4 can lead to good bounds for combinatorial configuration problems like those considered in [21].
@@ Acknowledgements
An earlier version of this manuscript presented the bound
$f(K)=O(\log(K)\log\log(K)\log\log\log(K))$
in Theorem 1.4. Fredy Yip (private communication) observed that Proposition 2.11 could be improved, with the eventual effect of removing the $\log\log\log(K)$ factor from this estimate. Special thanks to Fredy Yip for providing this optimization.
We would like to thank Terence Tao for a great deal of guidance and support. We would also like to thank James Leng for helpful conversations and encouragement, especially regarding the $U^{3}$ inverse theorem. Finally, we would like to thank Ben Green for helpful comments on a previous version of this manuscript. The author is partially supported by NSF Grant DMS-2347850.
@@ 2.1 Definitions and Notation
To simplify notation in this paper, we introduce the following convention regarding iterated logarithms.
@@
For a real number $x\geq 3$, we define $\log_{2}(x)=\log\log(x)$. In particular, we never use $\log_{2}$ to refer to the base $2$-logarithm.
@@
Entropy analogues of standard inequalities in additive combinatorics were developed by Ruzsa [22] and Tao [31] in 2009. Several authors have contributed to this program since then, developing an “entropic Ruzsa calculus” which allows us to perform many of the standard operations from additive combinatorics in the entropic setting. We record the estimates from this theory which we will use in Appendix A.
@@
Although the doubling constant and Ruzsa distance provide important information about the growth and dimension of a set $A\subseteq G$, they do not paint a complete picture of the approximate structure of $A$. We consider the following pair of standard examples, also discussed in [27] and [10].
@@
To distinguish between these instances of “covering” and “containment”, we define the following, which is analogous to a combinatorial polynomial growth condition studied by Sanders in [26], [27].
@@ 2.2 Strategy of Proof
Part 1: The Entropy Side. To motivate the first half of our argument, we will give a brief overview of the proof of the Polynomial Freiman-Ruzsa conjecture over $\mathbb{F}_{2}$ given by Gowers, Green, Manners, and Tao [10]. As in [10], we will obtain a structural characterization of the minimizers of a functional involving some entropic Ruzsa distances. More precisely, they prove the following.
Theorem 2.9 ([10], Propositions 2.1 and 2.2).
@@
Let $\eta=1/438$. For random variables $X,Y$ supported on $A$, we define
$\tau[X;Y]=\text{d}[X;Y]+\frac{\eta}{\log_{2}(3K)}(\text{d}[X;U_{A}]+\text{d}[Y;U_{A}]).$
@@
Let $C$ be a constant. Suppose $(X,Y)$ are minimizers of $\tau$. Then either $K=O_{C}(1)$, or $X$ has polynomial growth of order $d=O(\log(3K))$ at scale $Cd^{5}$. Moreover, $\text{d}[X;U_{A}]\leq 30d\log(d)$.
The constants 1/438 and 30 above can certainly be improved, but we make no effort to do so.
Part 2: The Fourier Side. The arguments in this part of the proof largely mirror those by Sanders in [27, Sections 9-11]. Sanders considered a combinatorial polynomial growth condition instead of an entropic one, and showed that a set with polynomial growth of a given order could be related to a convex coset progression with controlled size and dimension. More specifically, he showed the following.
Theorem 2.12 ([27], Theorem 2.7).
@@
The condition (2) is stronger than Definition 2.8 in the sense that if a set $X$ satisfies condition (2) with an exponent $d$, then $U_{X}$ has polynomial growth of order $d$ as well. The converse implication is not true; see [11, Appendix B] for an example of large gaps between combinatorial and entropic notions of doubling. However, we are able to show using modifications of Sanders’ arguments that under our weaker notion of polynomial growth, we may obtain a comparable amount of structure.
@@
In view of the quantitative similarities between Theorem 2.12 and Proposition 2.13, it is natural to wonder how our argument provides any quantitative improvement over the results in [27]. There, Sanders uses the Croot-Sisask lemma for almost-periods of convolutions [4] to produce a somewhat sparse set $T$ ($\text{d}[U_{T};U_{A}]\approx\log(K)^{3}$) with a relatively high order of polynomial growth ($d\approx\log(K)^{3}$). The information-theoretic part of our argument produces a much “denser” random variable with a much lower order of polynomial growth, which eventually leads to savings in the final bounds. The author would be very interested to see if the methods in this paper can be used to provide quantitative improvements to any of the many other applications of the Croot-Sisask lemma.
@@
Let $C$ be the constant from Proposition 2.13. We now select a minimizer $(X,Y)$ of $\tau$ and apply Proposition 2.11. If $K=O(1)$, we are done, since we can apply Theorem 1.3 to cover $A$ by $O(1)$ translates of a convex coset progression of size at most $O(|A|)$ and dimension at most $O(1)$. Otherwise, $X$ and $A$ satisfy the hypotheses of Proposition 2.13 with $d=O(\log(3K))$. There is thus a convex coset progression $M$ with dimension $O(d\log(d))$ and size at most $\exp(O(d\log(d))|A|$ such that $A$ can be covered by at most $K\exp(O(d\log(d)))$ translates of $M$. Since $d\log(d)=O(\log(3K)\log_{2}(3K))$, the claim follows. ∎
@@ 3 Obtaining a Random Variable with Polynomial Growth
Remark. In a previous version of this manuscript, Proposition 3.1 was stated using a slightly different penalty term in (3). Fredy Yip (private communication) observed that the penalty term could be optimized, leading ultimately to a slight improvement to Theorem 1.4.
@@
$\text{d}[X|nX;Y|nY]\leq\log(k)-\frac{73\eta\log(k)}{\log_{2}(3K)}(\log(n)-\log(n-1))\leq\log(k)-\frac{73\eta\log(k)}{n\log_{2}(3K)}.$
@@
$\frac{36}{n}\log(k)$
since $m\geq n/3$, $m\leq\log(K)^{6}$, and $6\cdot 73\eta\leq 1$.
By a similar argument, $\text{d}[Y|nY;U_{A}]-\text{d}[Y;U_{A}]\leq\frac{36}{n}\log(k)$.
@@
$\text{d}[X|nX;Y|nY]-\text{d}[X;Y]\leq-\frac{73\eta\log(k)}{n\log_{2}(3K)},$
$\frac{\eta}{\log_{2}(3K)}(\text{d}[X|nX;U_{A}]-\text{d}[X;U_{A}])\leq\frac{36\eta\log(k)}{n\log_{2}(3K)},\text{ and}$
$\frac{\eta}{\log_{2}(3K)}(\text{d}[Y|nY;U_{A}]-\text{d}[Y;U_{A}])\leq\frac{36\eta\log(k)}{n\log_{2}(3K)}.$
@@
$\sum_{s,t\in G}\mathbb{P}(nX=t,nY=s)\tau[X|nX=t;Y|nY=s]-\tau[X;Y]\leq-\frac{73\eta\log(k)}{n\log_{2}(3K)}+\frac{72\eta\log(k)}{n\log_{2}(3K)}<0.$
There thus exist $s,t\in G$ such that $\tau[X|nX=t;Y|nY=s]<\tau[X;Y]$, so $(X,Y)$ could not have been a minimizer of $\tau$. ∎
@@
By Proposition 3.1, either $K=O(1)$, or each term is bounded above by
$\log(k)+\frac{73\eta}{\log_{2}(3K)}\log(2^{j})\log(k)\leq\log(k)+\frac{6\cdot 73\eta\log_{2}(K)}{\log_{2}(3K)}\leq 2\log(k),$
since $2^{j}\leq\log(K)^{6}$ and $6\cdot 73\cdot\eta\leq 1$.
We therefore have
$\mathbb{H}(2^{\ell}X)-\mathbb{H}(X)\leq 12\ell\log(k).$
@@
$\mathbb{H}(nX)-\mathbb{H}(X)\leq\mathbb{H}(2^{\ell}X)-\mathbb{H}(X)\leq 12\ell\log(k)\leq 24\log(n)\log(k).$
@@
$\log(k)=\text{d}[X;Y]\leq\tau[X;Y]\leq\tau[U_{A};U_{A}]\leq 3\log(K)\leq 3\log(3K).$
Thus for any $n\leq\frac{1}{2}\log(K)^{6}$,
$\mathbb{H}(nX)-\mathbb{H}(X)\leq 72\log(n)\log(3K).$
Set $d=72\log(3K)$. Since $d^{5}=o(\log(K)^{6})$, either $K=O_{C}(1)$, or
@@
It remains to show that $\text{d}[X;U_{A}]\leq 30d\log(d)$. Since $(X,Y)$ minimizes $\tau$, we have
$\frac{\eta}{\log_{2}(3K)}\text{d}[X;U_{A}]\leq\tau[X;Y]\leq\tau[U_{A};U_{A}]\leq 3\text{d}[U_{A};U_{A}]\leq 3\log(3K),$
@@
$\text{d}[X;U_{A}]\leq 3\eta^{-1}\log(3K)\log_{2}(3K)\leq 1500\log(3K)\log_{2}(3K).$
This is at most $30d\log(d)$. ∎